### v1

# Distilling the Knowledge in Data Pruning

Abstract
With the increasing size of datasets used for training neural networks, data pruning becomes an attractive field of research. However, most current data pruning algorithms are limited in their ability to preserve accuracy compared to models trained on the full data, especially in high pruning regimes. In this paper we explore the application of data pruning while incorporating knowledge distillation (KD) when training on a pruned subset. That is, rather than relying solely on ground-truth labels, we also use the soft predictions from a teacher network pre-trained on the complete data. By integrating KD into training, we demonstrate significant improvement across datasets, pruning methods, and on all pruning fractions. We first establish a theoretical motivation for employing self-distillation to improve training on pruned data. Then, we empirically make a compelling and highly practical observation: using KD, simple random pruning is comparable or superior to sophisticated pruning methods across all pruning regimes. On ImageNet for example, we achieve superior accuracy despite training on a random subset of only 50% of the data. Additionally, we demonstrate a crucial connection between the pruning factor and the optimal knowledge distillation weight. This helps mitigate the impact of samples with noisy labels and low-quality images retained by typical pruning algorithms. Finally, we make an intriguing observation: when using lower pruning fractions, larger teachers lead to accuracy degradation, while surprisingly, employing teachers with a smaller capacity than the student’s may improve results. Our code will be made available.
Keywords:
Data pruning Knowledge distillation Image classification
1 Introduction
(a) In knowledge distillation for model compression (left), a large teacher network is used to guide the training of a smaller student network. In contrast, here we investigate the usage of a teacher model, pre-trained on a full dataset, to guide a student model during training on a pruned subset of the data (right).
(b) Accuracy vs. pruning methods (CIFAR-100)
(c) Impact of teacher size (CIFAR-100)
Figure 1: Knowledge distillation for data pruning. (a) The difference between KD for model compression and KD for data pruning. (b) We find that by integrating KD into the training, simple random pruning outperforms other sophisticated pruning algorithms across all pruning regimes. (c) Interestingly, we observe that when using small data fractions, training with large teachers degrades accuracy, while smaller teachers are favored. This suggests that in high pruning regimes (low $f$), the training is more sensitive to the capacity gap between the teacher and the student.
Recently, data pruning has gained increased interest in the literature due to the growing size of datasets used for training neural networks. Algorithms for data pruning aim to retain the most representative samples of a given dataset and enable the conservation of memory and reduction of computational costs by allowing training on a compact and small subset of the original data. For instance, data pruning can be useful for accelerating hyper-parameter optimization or neural architecture search (NAS) efforts. It may also be used in continual learning or active learning applications.
Existing methods for data pruning have shown remarkable success in achieving good accuracy while retaining only a fraction, $f<1$, of the original data; see for example [38, 32, 10, 25] and the overview in [12]. However, those approaches are still limited in their ability to match the accuracy levels obtained by models trained on the complete dataset, especially in high compression regimes (low $f$).
Score-based data pruning algorithms typically rely on the entire data to train neural networks for selecting the most representative samples. The ‘forgetting’ method [38] counts for each sample the number of instances during training where the network’s prediction for that sample shifts from “correct” to “misclassified”. Samples with high rates of forgetting events are assigned higher scores as they are considered harder and more valuable for the training. The GraNd and EL2N methods [32] compute a score for each sample based on the gradient norm (GraNd) or the error L2-norm (EL2N) between the network’s prediction and the ground-truth label, respectively. The scores are computed and averaged over an ensemble of models trained on the full dataset. For each method, we note that once the sample scores are calculated, the models trained on the full dataset are discarded and are no longer in use.
In this paper, we explore the benefit of using a model trained on a complete dataset to enhance training on a pruned subset of the data using knowledge distillation (KD). The motivation behind this approach is that a teacher model trained on the complete dataset captures essential information and core statistics about the entire data. This knowledge can then be utilized when training on a pruned subset. While KD has been extensively studied and demonstrated significant improvements in tasks such as model compression, herein we aim to investigate its impact in the context of data pruning and propose innovative findings for practical usage. Note that, in contrast to traditional model compression techniques, our approach focuses on self-distillation (SD), where the teacher and student have identical architectures. An overview of the proposed training scheme is presented in Fig. 0(a).
We experimentally demonstrate that incorporating the (soft) predictions provided by the teacher throughout the training process on the pruned data significantly and consistently improves accuracy across multiple datasets, various pruning algorithms, and all pruning fractions (see Fig. 0(b) for example). In particular, using KD, we can achieve comparable or even higher accuracy with only a small portion of the data (e.g., retaining $50\%$ and $10\%$ of the data for CIFAR-100 and SVHN, respectively). Moreover, a dramatic improvement is achieved especially for small pruning fractions (low $f$). For example, on CIFAR-100 with pruning factor $f=0.1$, accuracy improves by $17\%$ (from $39.8\%$ to $56.8\%$) using random pruning. On ImageNet with $f=0.1$, the Top-5 accuracy increases by $5\%$ (from $82.37\%$ to $87.19\%$) using random pruning, and by $20\%$ (from $62.47\%$ to $82.47\%$) using EL2N. To explain these improvements, we provide theoretical motivation for integrating SD when training on pruned data. Specifically, we show that using a teacher trained on the entire data reduces the bias of the student’s estimation error.
In addition, we present several empirical key observations. First, our results demonstrate that simple random pruning outperforms other sophisticated pruning algorithms in high pruning regimes (low $f$), both with and without knowledge distillation. Notably, prior research demonstrated this phenomenon in the absence of KD [35, 48]. Second, we demonstrate a useful connection between the pruning factor $f$ and the optimal weight of the KD loss. Generally, utilizing data pruning algorithms to select high-scoring samples amplifies sensitivity to samples with noisy labels or low quality. This is because keeping the hardest samples increases the portion of these samples as we retain a smaller data fraction. Based on this observation, we propose to adapt the weight of the KD loss according to the pruning factor. That is, for low pruning factors, we should increase the contribution of the KD term as the teacher’s soft predictions reflect possible label ambiguity embedded in the class confidences. On the other hand, when the pruning factor is high, we can decrease the contribution of the KD term to rely more on the ground-truth labels.
Finally, we observe a striking phenomenon when training with KD using larger teachers: in high pruning regimes (low $f$), the optimization becomes significantly more sensitive to the capacity gap between the teacher and the student model. This relates to the well known capacity gap problem [26]. Interestingly, we find that for small pruning fractions, the student benefits more from teachers with equal or even smaller capacities than its own, see Fig. 0(c).
The contributions of the paper can be summarized as follows:
  • We explore the utilization of knowledge distillation while training on pruned data by leveraging the soft predictions of a teacher pre-trained on the complete data. Using KD, we demonstrate consistent improvement across all pruning regimes, various pruning algorithms, and on multiple datasets.
  • We theoretically show, for the case of linear regression, that using a teacher trained on the entire data reduces the bias of the student’s estimation error.
  • We find that KD is robust to the choice of the pruning mechanism for high pruning fractions. In particular, using KD, random pruning achieves comparable or superior accuracy compared to other sophisticated methods across all pruning regimes.
  • We demonstrate that by appropriately choosing the KD weight, one can mitigate the impact of label noise and low-quality samples that are retained by common pruning algorithms.
  • We make the striking observation that, for small pruning fractions, increasing the teacher size degrades accuracy, while, intriguingly, using teachers with smaller capacities than the student’s improves results.
2 Related work
Data pruning. Data pruning, also known as coreset selection [27, 18, 37], refers to methods aiming to reduce the dataset size for training neural networks. Recent approaches have shown significant progress in retaining less data while maintaining high classification accuracy [38, 32, 10, 25, 5, 35]. In [35], the authors showed theoretically and empirically that data pruning can improve the power law scaling of the dataset size by choosing an optimal pruning fraction as a function of the initial dataset size. Additionally, studies in [35, 2] have demonstrated that existing pruning algorithms often underperform when compared to random pruning methods, especially in high pruning regimes. Recently, in [48], the authors suggested a theoretical explanation to this accuracy drop, and proposed a coverage-centric pruning approach which better handles the data coverage.
Data pruning proves valuable at reducing memory and computational cost in various applications, including tasks such as hyper-parameter search [6], neural architecture search [8], continual and incremental learning [22], as well as active learning [27, 5].
Other related fields are dataset distillation and data-free knowledge distillation (DFKD). Dataset distillation approaches [40, 47, 45] aim to compress a given dataset by synthesizing a small number of samples from the original data. The goal of DFKD is to employ model compression in scenarios where the original dataset is inaccessible, for example, due to privacy concerns. Common approaches for DFKD involve generating synthetic samples suitable for KD [24, 44] or inverting the teacher’s information to reconstruct synthetic inputs [28, 42]. Recently, the works in [7, 43], utilized pseudo labels in training with dataset distillation. Unlike dataset distillation and DFKD, which include synthetic data generation, our work focuses on enhancing models trained on pruned datasets created through sample selection, using KD. Moreover, in this paper we present a collection of innovative and practical findings for the application of KD in data pruning.
Knowledge distillation. Knowledge distillation is a popular method aiming at distilling the knowledge from one network to another. It is often used to improve the accuracy of a small model using the guidance of a large teacher network [3, 15]. In recent years, numerous variants and extensions of KD have been developed. For example, [46, 33] utilized feature activations from intermediate layers to transfer knowledge across different representation levels. Other methods have proposed variants of KD criteria [41, 17, 19, 1], as well as designing objectives for representation distillation, as demonstrated in [36, 4]. Self-distillation (SD) refers to the case where the teacher and student have identical architectures. It has been demonstrated that accuracy improvement can be achieved using SD [11]. Recently, theoretical findings were introduced for self-distillation in the presence of label noise [9].
In our paper, we explore the process of distilling knowledge from a model trained on a large dataset to a model trained on a pruned subset of the original data. Specifically, we focus on self-distillation and present several striking observations that emerge when integrating SD for data pruning.
3 Method
Given a dataset $\mathcal{D}$ with $N$ labeled samples $\{x_{i},y_{i}\}_{i=1}^{N}$, a data pruning algorithm $\mathcal{A}$ aims at selecting a subset $\mathcal{P}\subset\mathcal{D}$ of the most representative samples for training. We denote by $f$ the pruning factor, which represents the fraction of data to retain, calculated as $f=N_{f}/N$ where $N_{f}$ is the size of the pruned dataset. Note that $10<f<$ 1. Score-based algorithms assign a score to each sample, representing its importance in the learning process. Let $s_{i}$ be the score corresponding to a sample $x_{i}$, sorting them in a descending order $s_{k_{1}}>s_{k_{2}},...,>s_{k_{N}}$, following the sorting indices $\{k_{1},...,k_{N}\}$, we obtain the pruned dataset by retaining the highest scoring samples, $\mathcal{P}=\{x_{k_{1}},...,x_{k_{N_{f}}}\}$. Usually, score-based algorithms retain hard samples while excluding the easy ones. Note that in random pruning, we simply sample the indices $k_{1},...,k_{N}$ uniformly. In this paper, given a pruning algorithm $\mathcal{A}$, our objective is to train a model on the pruned dataset $\mathcal{P}$ while maximizing accuracy.
3.1 Training on the pruned dataset using KD
Typically, score-based pruning methods involve training multiple models on the full dataset $\mathcal{D}$ to compute the scores [38, 32, 10, 25]. These models are discarded and are not utilized further after the scores are computed. We argue that a model trained on the full dataset encapsulates valuable information about the entire distribution of the data and its classification boundaries, which can be leveraged when training on the pruned data $\mathcal{P}$. In this work, we investigate a training scheme which incorporates the soft predictions of a teacher network, pre-trained on the full dataset, throughout training on the pruned data.
Let $f_{t}(x)$ be the teacher backbone pre-trained on $\mathcal{D}$. The teacher outputs logits $\{z_{i}\}_{i=1}^{C}$, where $C$ is the number of classes. The teacher’s soft predictions are computed by,
$q_{i}=\frac{\exp(z_{i}/\tau)}{\sum_{j}\exp(z_{j}/\tau)},\quad i=1\dots C,$ (1)
where $\tau$ is the temperature hyper-parameter. Similarly, we denote the student model trained on the dataset $\mathcal{P}$ as $f_{s}(x;\mathbf{\theta})$, where $\mathbf{\theta}$ represents the student’s parameters. The student’s $i$-th soft prediction is denoted by $p_{i}(\mathbf{\theta})$. We optimize the student model using the following loss function,
$\mathcal{L}(\theta)=(1-\alpha)\mathcal{L}_{\text{cls}}(\theta)+\alpha\mathcal{L}_{\text{KD}}(\theta),$ (2)
where the classification loss $\mathcal{L}_{\text{cls}}(\theta)$ measures the cross-entropy between the ground-truth labels and the student’s predictions, represented as: $-\sum_{i}y_{i}\log p_{i}(\theta)$. For the KD term $\mathcal{L}_{\text{KD}}(\theta)$, a common choice is the Kullback-Leibler (KL) divergence between the soft predictions of the teacher and the student. The hyper-parameter $\alpha$ controls the weight of the KD term relative to the classification loss.
Integrating the KD loss into the training process allows us to leverage the valuable knowledge embedded in the teacher’s soft predictions $q_{i}$. These predictions may encapsulate potential relationships between categories and class hierarchies, accumulated by the teacher during its training on the entire dataset. To illustrate this, we provide a qualitative example in Fig. 2 that presents the soft predictions generated for a specific sample from the CIFAR-100 dataset. CIFAR-100 comprises 100 classes, organized into 20 super-classes, each containing 5 sub-classes. For example, the super-class "People" contains the classes: "Baby", "Boy", "Girl", "Man", and "Woman". As shown in Fig. 2 (top), the teacher accurately predicts the ground-truth class "Girl" (class index 35) with high confidence while also assigning high confidence values to the classes "Woman" (98), "Man" (46), and "Boy" (11). This ‘dark knowledge’ is valuable for training as it offers a broader view of class hierarchies and data distribution. Fig. 2 (middle) illustrates that a model trained on only $25\%$ of the data fails to capture such class relationships. Intuitively, reliable data and class distributions can be effectively learned from large datasets, but are harder to infer from small datasets. Conversely, in Fig. 2 (bottom) we show that using knowledge distillation, the student successfully learns these delicate data relationships from the teacher despite training only on the pruned data.
Figure 2: Learning from the teacher predictions. An example of soft predictions computed by a teacher model trained on the entire data (top), a model trained on $25\%$ of the data (middle), and a student model trained on $25\%$ of the data with KD (bottom), for an evaluation sample of class "Girl" from CIFAR-100. Using KD, the student can better learn close or ambiguous categories by leveraging knowledge captured by the teacher from the full dataset.
In Sec. 4.1, we empirically demonstrate that integrating knowledge distillation into the optimization process of the student model, trained on pruned data, leads to significant improvements across all pruning factors and various pruning methods. In addition, we show that simple random pruning outperforms other sophisticated pruning methods for low pruning fractions (low $f$), both with and without knowledge distillation. We note that prior work has demonstrated this phenomenon in the absence of KD [35]. Interestingly, we also observe that training with KD is robust to the choice of the data pruning method, including simple random pruning, for sufficiently high pruning fractions.
These observations on the effectiveness of random pruning in the presence of KD are compelling, especially in scenarios where data pruning occurs unintentionally as a by-product of the system, such as cases where the full dataset is no longer accessible due to privacy concerns. However, using knowledge distillation we can train a student model on the remaining available data while maintaining a high level of accuracy.
3.2 Mitigating noisy samples in pruned datasets
In general, hard samples are essential for the optimization process as they are located close to the classification boundaries. However, retaining the hardest samples while excluding moderate and easy ones increases the proportion of samples with noisy and ambiguous labels, or images with poor quality. For example, in Fig. 3, we present the highest scoring images selected by the ‘forgetting’ pruning algorithm for CIFAR-100 and SVHN. As can be seen, in the majority of the images determining the class is non-trivial due to the complexity of the category (e.g., fine-grained classes) or due to poor quality. By using knowledge distillation the student can learn such label ambiguity and mitigate noisy labels.
In a recent work [9] it was demonstrated that the benefit of using a teacher’s predictions increases with the degree of label noise. Consequently, it was found that more weight should be assigned to the KD term as the noise variance increases. Similarly, in our work we empirically demonstrate that as the pruning factor $f$ becomes lower, we should rely more on the teacher’s predictions by increasing $\alpha$ in Eq. 2. Conversely, as the pruning factor is increased, we may rely more on the ground-truth labels by decreasing $\alpha$. We find that setting $\alpha$ properly is crucial when applying pruning methods that retain hard samples. Formally, the objective should be aware of the pruning fraction $f$ as follows,
$\mathcal{L}(\theta,f)=\bigl{(}1-\alpha(f)\bigr{)}\mathcal{L}_{\text{cls}}(\theta)+\alpha(f)\mathcal{L}_{\text{KD}}(\theta).$ (3)
For example, as can be seen from Fig. 6, when the pruning fraction is low ($f=0.1$), training with $\alpha=1$ is superior, achieving more than $8\%$ higher accuracy compared to $\alpha=0.5$. Conversely, for high pruning fractions (e.g. $f=0.7$), using $\alpha=0.5$ outperforms $\alpha=1$ by more than $1\%$ accuracy. We further explore the relationship between $\alpha$ and $f$ in Sec. 4.2.
Refer to caption (a) CIFAR-100 highest score pruning samples Refer to caption (b) SVHN highest score pruning samples
Figure 3: Highest scoring samples. Top 10 highest scoring samples selected by the ‘forgetting’ method for CIFAR-100 and SVHN datasets. The labels of the majority of the images are ambiguous or hard to resolve due to class complexity or low image quality.
3.3 Theoretical motivation
In this section we provide a theoretical motivation for the success of self-distillation in enhancing training on pruned data. We base our analysis on the recent results reported in [9] for the case of regularized linear regression. Note that while we use logistic regression in practice, we anchor our theoretical results in linear regression for the sake of simplicity. Also, it often allows for a reliable emulation of outcomes observed in processes applied to logistic regression (see e.g. in  [9]). In particular, we show that employing self-distillation using a teacher model trained on a larger dataset reduces the error bias of the student estimation.
We are given a data matrix, $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{N}]\in\mathbb{R}^{d\times N}$, and a corresponding label vector $\mathbf{y}=[y_{1},\dots,y_{N}]\in\mathbb{R}^{N}$, where $N$ and $d$ are the number of samples and their dimension, respectively. Let $\boldsymbol{\theta}^{*}\in\mathbb{R}^{d}$ be the ground-truth model parameters. The labels are assumed to be random variables, linearly modeled by $\mathbf{y}=\mathbf{X}^{T}\boldsymbol{\theta}^{*}+\boldsymbol{\eta}$, where $\boldsymbol{\eta}\in\mathbb{R}^{N}$ is assumed to be Gaussian noise, uncorrelated and independent on the observations. In data pruning, we select $N_{f}$ columns from $\mathbf{X}$ and their corresponding labels: $\mathbf{X}_{f}\in\mathbb{R}^{d\times N_{f}}$, $\mathbf{y}_{f}\in\mathbb{R}^{N_{f}}$. Thus, $\mathbf{y}_{f}=\mathbf{X}_{f}^{T}\boldsymbol{\theta}^{*}+\boldsymbol{\eta}_{f}$. We also assume that $d\leq N_{f}\leq N$ which is true in most practical scenarios. Solving linear regularized regression using pruned dataset with fraction $f$, the parameters are obtained by:
$\hat{\boldsymbol{\theta}}(f)$ $=\arg\!\min_{\boldsymbol{\theta}}\biggl{\{}||\mathbf{y}_{f}-\mathbf{X}_{f}^{T}\mathbf{\boldsymbol{\theta}}||_{2}^{2}+\frac{\lambda}{2}||\boldsymbol{\theta}||_{2}^{2}\biggr{\}}$
$=(\mathbf{X}_{f}\mathbf{X}_{f}^{T}+\lambda\mathbf{I}_{d})^{-1}\mathbf{X}_{f}\mathbf{y}_{f},$
where $\lambda>0$ is the regularization hyper-parameter, and $\mathbf{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix. Note that a teacher trained on the full data is given by: $\hat{\boldsymbol{\theta}}_{t}=\hat{\boldsymbol{\theta}}(1)=(\mathbf{X}\mathbf{X}^{T}+\lambda\mathbf{I}_{d})^{-1}\mathbf{X}\mathbf{y}$.
Here, we look at the more general case where the student is trained on a pruned subset with factor $f$, and the teacher model is trained on a larger subset of the data, $f_{t}>f$. Following  [9], the model learned by the student is given by,
$\hat{\boldsymbol{\theta}}_{s}(\alpha,f,f_{t})=(1-\alpha)(\mathbf{X}_{f}\mathbf{X}_{f}^{T}+\lambda\mathbf{I}_{d})^{-1}\mathbf{X}_{f}\mathbf{y}_{f}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\alpha(\mathbf{X}_{f}\mathbf{X}_{f}^{T}+\lambda\mathbf{I}_{d})^{-1}\mathbf{X}_{f}\hat{\mathbf{y}}_{f}^{(t)}$ (4)
$=(\mathbf{X}_{f}\mathbf{X}_{f}^{T}+\lambda\mathbf{I}_{d})^{-1}\mathbf{X}_{f}\bigl{(}(1-\alpha)\mathbf{y}_{f}+\alpha\mathbf{X}_{f}^{T}\hat{\boldsymbol{\theta}}(f_{t})\bigr{)},$
where $\hat{\mathbf{y}}_{f}^{(t)}=\mathbf{X}_{f}^{T}\hat{\boldsymbol{\theta}}(f_{t})$, i.e., the teacher’s predictions of the student’s samples $\mathbf{X}_{f}$. Note that in a regular self-distillation (without pruning), we have $f=f_{t}=1$, and $\alpha>0$. Also, in a regular training on pruned data (without KD), $f<1$, and $\alpha=0$. In our scenario we utilize self-distillation for data pruning, i.e., $f<f_{t}\leq 1$, and $\alpha>0$.
We denote the student estimation error as $\boldsymbol{\epsilon}_{s}(\alpha,f,f_{t})=\hat{\boldsymbol{\theta}}_{s}(\alpha,f,f_{t})-\boldsymbol{\theta}^{*}$. In  [9], the authors show that employing self-distillation ($\alpha>0)$ reduces the variance of the student estimation, but on the other hand, increases its bias. In the following, we show that distilling the knowledge from a teacher trained on a larger data subset w.r.t the student, decreases the error estimation bias.
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{d\times N}$ and $\mathbf{y}\in\mathbb{R}^{N}$ be the full observation matrix and label vector, respectively. Let $\mathbf{y}_{f}=\mathbf{X}_{f}^{T}\boldsymbol{\theta}^{*}+\boldsymbol{\eta}_{f}$, where $\boldsymbol{\theta}^{*}$ is the ground-truth projection vector and $\boldsymbol{\eta}_{f}\in\mathbb{R}^{N}$ is a Gaussian uncorrelated noise independent on $\mathbf{X}$. Let $\boldsymbol{\epsilon}_{s}(\alpha,f,f_{t})=\hat{\boldsymbol{\theta}}_{s}(\alpha,f,f_{t})-\boldsymbol{\theta}^{*}$ be the student estimation error. Also, assume that $d\leq N_{f}\leq N$, and $f\leq f_{t}$. Then, for any $\alpha$,
$||\mathbb{E}_{\eta}[\boldsymbol{\epsilon}_{s}(\alpha,f,f_{t})]||^{2}\leq||\mathbb{E}_{\eta}[\boldsymbol{\epsilon}_{s}(\alpha,f,f))]||^{2}.$
We include the proof for Theorem  1 in the supplementary. As data pruning is susceptible to label noise due to retaining the hardest samples, this finding demonstrates the utility of the proposed method. It suggests that employing self-distillation with a teacher trained on the entire dataset ($f_{t}=1)$ enables the reduction of estimation bias in a student trained on a pruned subset.
4 Experimental results
In this section we provide empirical evidence for our method through extensive experimentation over a variety of datasets, an assortment of data pruning methods and across a wide range of pruning levels. Then, we also investigate how the KD weight, the teacher size and the KD method affect student performance under different pruning regimes.
(a) CIFAR-100
(b) SVHN
(c) CIFAR-10
Figure 4: Data pruning results with knowledge distillation. Accuracy results across different pruning factors $f$, and various pruning approaches (’forgetting’, EL2N, GraNd and random pruning) on the CIFAR-100, SVHN, and CIFAR-10 datasets. We use an equalized weight in the loss (i.e., $\alpha=0.5$). Using KD, significant improvement is achieved across all pruning regimes and all pruning methods. Random pruning outperforms other pruning methods for low pruning factors. For sufficiently high $f$, the accuracy is robust to the choice of the pruning approach in the presence of KD.
(a) ImageNet, Top-1 accuracy
(b) ImageNet, Top-5 accuracy
Figure 5: Data pruning results with KD on ImageNet. Accuracy results across different pruning factors $f$, and various pruning methods on the ImageNet dataset. We use an equalized weight ($\alpha=0.5$) in Eq. 2.
Datasets. We perform experiments on four classification datasets: CIFAR-10 [20] with 10 classes, consists of 50,000 training samples and 10,000 testing samples; SVHN [29] with 10 classes, consists of 73,257 training samples and 26,032 testing samples; CIFAR-100 [21] with 100 classes, consists of 50,000 training samples and 10,000 testing samples; and ImageNet [34] with 1,000 classes, consists of 1.2M training samples and 50K testing samples.
Pruning Methods. We utilize several score-based data-pruning algorithms: ‘forgetting’ [38], Gradient Norm (GraNd), Error L2-Norm (EL2N) [32] and ‘memorization’ [^1] We note that while the authors of memorization did not originally utilize the method for data pruning, its efficacy on ImageNet was later demonstrated by [35]. [10]. We also utilize a class-balanced random pruning scheme, which, given a pruning budget, randomly and equally draws samples from each class.
Implementation Details. For computational efficiency we conduct our self-distillation experiments on all datasets using the ResNet-32 [13] architecture, except for ImageNet for which we utilize the larger ResNet-50. Our training and distillation recipes are simple. We utilize SGD with Momentum to optimize the models and incorporate basic data-augmentations during training. Additional implementation details can be found in the supplementary.
4.1 Training on pruned data with KD
To demonstrate the advantage of incorporating KD-based supervision when training on pruned data, we utilize the aforementioned data pruning methods on each dataset using a wide range of pruning factors. Then, we train models on the produced data subsets with and without KD. We note that in the presence of KD the respective teachers that are utilized are trained on the full datasets.
As can be observed in Figs. 4 and 5, the incorporation of KD into the training process consistently enhances model accuracy across all of the tested scenarios, regardless of the tested dataset, pruning method or pruning level. For example, compared to baseline models trained on the full datasets without KD, utilizing KD can lead to comparable accuracy levels by retaining only small portions of the original datasets (e.g., 10%, 30%, 50% on SVHN, CIFAR-10, and CIFAR-100, respectively, using ‘forgetting’). In fact, even on a large scale dataset as ImageNet, comparable accuracy can be achieved by randomly retaining just 30% of the data, while training on larger subsets remarkably results in superior accuracy to the baseline (e.g., +1.6% using a random subset of 70%). In contrast, we observe that in the absence of KD, preserving accuracy is only possible in the presence of more data (i.e., more relaxed pruning levels). In particular, on CIFAR-100 we found accuracy preservation to be unattainable without KD even in the least aggressive pruning scenario we have tested ($f=0.9$), and regardless of the pruning method.
Figure 6: Optimal KD weight versus pruning factor. Accuracy is presented for CIFAR-100 while varying the KD weight $\alpha$ for different pruning factors. We utilize ‘forgetting’ as the pruning method. For low pruning fractions (low $f$), accuracy generally increases when increasing the KD weight to rely more on the teacher’s soft predictions. As we use higher pruning fractions (high $f$), it is usually better to lower $\alpha$ in order to increase the contribution of the ground-truth labels.
Moreover, we note that the accuracy gains due to KD are most significant in high-compression scenarios. For instance, on CIFAR-100 with $f=0.1$, KD contributes to absolute accuracy improvements of 17%, 22.4%, 21%, and 19.7% across the random, ‘forgetting’, GraNd, and EL2N pruning methods, respectively. Similarly, on SVHN, which permits even stronger compression, improvements of the same order of magnitude can be observed at a lower pruning factor ($f=0.01$).
These findings support the idea that the soft-predictions produced by a well-informed teacher contain rich and valuable information that can greatly benefit a student in a limited-data setting. This ‘dark knowledge’, notably absent in conventional one-hot labels, allows the student to deduce stronger generalizations from each available data sample, which in turn translates to better performance given the same training data.
Finally, two additional interesting patterns emerge from our experiments. First, in high-compression scenarios (e.g., $f\leq 0.4$ in CIFAR-100, $f\leq 0.08$ in SVHN), it is evident that random pruning surpasses all other methods in effectiveness, both with and without KD. This aligns with the notion that aggressive pruning via score-based techniques retains larger concentrations of low quality or noisy samples due to mistaking them for challenging cases. This phenomenon was previously noted without KD in [35]. Second, under low-compression conditions (e.g., $f\geq 0.5$ in CIFAR-100, $f\geq 0.2$ in SVHN), we observe that KD renders the student model robust to the pruning technique used. This finding is significant as it suggests that it may be possible to forgo state-of-the-art pruning techniques in favor of basic random pruning in the presence of KD.
4.2 Adapting the KD weight vs. the pruning factor
We wish to investigate how varying the KD weight $\alpha$ affects the performance of the student under different pruning levels of a given dataset. To explore this we conduct experiments on CIFAR-100 with ’forgetting’ as the pruning method and present the results in Fig. 6. As can be observed, lower pruning fractions favor higher values of $\alpha$, while higher pruning fractions advocate for lower ones. As explained earlier, aggressive pruning via score-based methods tends to result in subsets with greater proportions of label noise and low quality samples. Hence, for lower pruning factors, increasing the KD weight seems to help the student mitigate the extra noise by relying more on the teacher’s predictions. Conversely, as the pruning factor increases and the proportions of noise in the pruned subset gradually diminish, it appears to be beneficial for the student to balance the contributions of KD and the ground-truth labels. Similar results on SVHN can be found in the supplementary.
(a) CIFAR-100. Student architecture: ResNet-32 (RN32).
(b) CIFAR-100. Student architecture: ResNet-20 (RN20).
(c) CIFAR-10. Student architecture: ResNet-32 (RN32).
Figure 7: Exploring the effect of the teacher’s capacity. Accuracy results for a student with (a) ResNet-32 and (b) ResNet-20 architectures while using teacher models with increasing capacities along the horizontal axes. In each instance, we denote the teacher whose architecture matches that of the student by ‘SD’ (self-distillation). We use random pruning with different fractions. Surprisingly, under low pruning factors, increasing the teacher’s capacity results in lower student accuracy.
4.3 Using teachers of different capacities
Until now, we have focused on the case where both the student and teacher share the same architecture (i.e., self-distillation). In this section, we explore how the capacity of the teacher affects the student’s performance across different pruning regimes. In Fig. 6(a), we present accuracy results across various pruning factors for the case of randomly pruning CIFAR-100 and training with a ResNet-32 student. We employ 6 teacher architectures of increasing capacities: (1) ResNet-14 with $69.9\%$ accuracy, (2) ResNet-20 with $70.23\%$ accuracy, (3) ResNet-32 with $71.6\%$ accuracy, (4) ResNet-56 with $72.7\%$ accuracy, (5) ResNet-110 with $74.4\%$ accuracy, and (6) WRN-40-2 with $75.9\%$ accuracy. Also, note that for each teacher architecture we experiment with five different temperature values in the range $2-7$. We show the impact of the temperature selection in the supplementary. Similarly, in Fig. 6(b) we present results for the same experiment using a ResNet-20 student, while Fig. 6(c) depicts results of a similar experiment on CIFAR-10 for the ResNet-32 student.
As observed, at low pruning factors, increasing the teacher’s capacity harms the accuracy of the student. Conversely, at high pruning factors (e.g. $0.9$ and $1$) larger teachers generally lead to improved results, as expected in common KD for model compression. This trend is consistently observed across various student architectures and datasets, and is robust to the selection of the KD temperature. Additional results are provided in the supplementary.
This observation highlights a striking phenomenon: the capacity gap problem, which denotes the disparity in architecture size between the teacher and student, becomes more pronounced when applying knowledge distillation during training on pruned data. Surprisingly, for small pruning fractions, teachers with a smaller capacity than the student can lead to higher student accuracy levels compared to larger teachers. The results also suggest that the effectiveness of smaller teachers in improving the student is bounded to an extent by their own accuracy levels.
4.4 Comparing different KD approaches
Method 5% 10% 30% 50%
w/o KD
14.46
22.21
49.41
67.47
KD [15]
28.62
46.27
66.82
70.95
FitNets [33]
25.66
44.84
65.7
70.77
AB [14]
30.5
47.68
66.15
71.22
AT [46]
28.26
42.59
65.75
70.45
FT [19]
28.34
44.01
64.95
70.75
FSP [41]
27.62
37.16
62.79
69.72
NST [17]
26.2
44.5
64.93
70.97
PKT [31]
27.3
44.09
65.22
70.7
RKD [30]
21.69
43.03
65.43
70.36
SP [39]
29.09
42.53
65.62
70.72
VID [1] 32.5 49.46 67.38 71.16
Table 1: Comparison of different KD approaches on several pruning levels of CIFAR-100. We add various KD loss terms to Eq. 2, in addition to the vanilla KD term. ‘Forgetting’ is utilized as the pruning method. As observed, integrating VID [1] further improves training on the pruned dataset.
So far, we have utilized solely vanilla KD during training. Next we explore integrating additional KD approaches to the loss. In particular, we add an additional KD loss term $\mathcal{L}_{R}$ as follows: $\mathcal{L}(\theta)=\mathcal{L}_{\text{cls}}(\theta)+\alpha\mathcal{L}_{\text{KD}}(\theta)+\beta\mathcal{L}_{R}(\theta)$, where $\beta$ is a hyper-parameter. In this experiments, we simply set $\alpha$ and $\beta$ to $1$. In Tab. 1 we compare the performance of different KD methods on CIFAR-100 under low and average compression regimes. For a fair comparison, for the case of employing only the vanilla KD, we set $\alpha=2$, and $\beta=0$. As can be observed, integrating the Variational Information Distillation (VID) loss [1] improves results considerably for the tested cases. These results suggest that further improvement can be achieved by incorporating additional approaches to extract knowledge from the teacher.
5 Conclusion
In this paper, we investigated the application of knowledge distillation for training models on pruned data. We demonstrated the significant benefits of incorporating the teacher’s soft predictions into the training of the student across all pruning fractions, various pruning algorithms and multiple datasets. We empirically found that incorporating KD while using simple random pruning can achieve comparable or superior accuracy compared to sophisticated pruning approaches. We also demonstrated a useful connection between the pruning factor and the KD weight, and propose to adapt $\alpha$ accordingly. Finally, for small pruning fractions, we made the surprising observation that the student benefits more from teachers with equal or even smaller capacities than that of its own, over teachers with larger capacities.
We believe that these empirical observations can pave the way for in-depth research, both theoretical and experimental, to enhance our understanding of the optimization process when training models on pruned data using KD.

### v2

@@
 Our code will be made available.
-Keywords:
-Data pruning Knowledge distillation Image classification
 1 Introduction
-(a) In knowledge distillation for model compression (left), a large teacher network is used to guide the training of a smaller student network.
-In contrast, here we investigate the usage of a teacher model, pre-trained on a full dataset, to guide a student model during training on a pruned subset of the data (right).
+(a) KD for data pruning
 (b) Accuracy vs. pruning methods (CIFAR-100)
@@
 Figure 1: Knowledge distillation for data pruning.
-(a) The difference between KD for model compression and KD for data pruning.
+(a) We investigate the usage of a teacher model, pre-trained on a full dataset, to guide a student model during training on a pruned subset of the same data.
 (b) We find that by integrating KD into the training, simple random pruning outperforms other sophisticated pruning algorithms across all pruning regimes.
@@
 While KD has been extensively studied and demonstrated significant improvements in tasks such as model compression, herein we aim to investigate its impact in the context of data pruning and propose innovative findings for practical usage.
-Note that, in contrast to traditional model compression techniques, our approach focuses on self-distillation (SD), where the teacher and student have identical architectures.
-An overview of the proposed training scheme is presented in Fig.
-0(a).
+Note that, in contrast to traditional model compression techniques, here we focus on self-distillation (SD), where the teacher and student have identical architectures.
+The training scheme is illustrated in Fig.
+1(a).
 We experimentally demonstrate that incorporating the (soft) predictions provided by the teacher throughout the training process on the pruned data significantly and consistently improves accuracy across multiple datasets, various pruning algorithms, and all pruning fractions (see Fig.
-0(b) for example).
+1(b) for example).
 In particular, using KD, we can achieve comparable or even higher accuracy with only a small portion of the data (e.g., retaining $50\%$ and $10\%$ of the data for CIFAR-100 and SVHN, respectively).
@@
 Interestingly, we find that for small pruning fractions, the student benefits more from teachers with equal or even smaller capacities than its own, see Fig.
-0(c).
+1(c).
 The contributions of the paper can be summarized as follows:
-  • We explore the utilization of knowledge distillation while training on pruned data by leveraging the soft predictions of a teacher pre-trained on the complete data.
-Using KD, we demonstrate consistent improvement across all pruning regimes, various pruning algorithms, and on multiple datasets.
+  • Utilizing KD in data pruning, we find that training is robust to the choice of pruning mechanism at high pruning fractions.
+Notably, random pruning with KD achieves comparable or superior accuracy compared to other sophisticated methods across all pruning regimes.
   • We theoretically show, for the case of linear regression, that using a teacher trained on the entire data reduces the bias of the student’s estimation error.
-  • We find that KD is robust to the choice of the pruning mechanism for high pruning fractions.
-In particular, using KD, random pruning achieves comparable or superior accuracy compared to other sophisticated methods across all pruning regimes.
   • We demonstrate that by appropriately choosing the KD weight, one can mitigate the impact of label noise and low-quality samples that are retained by common pruning algorithms.
@@
 Recently, in [49], the authors suggested a theoretical explanation to this accuracy drop, and proposed a coverage-centric pruning approach which better handles the data coverage.
-Data pruning proves valuable at reducing memory and computational cost in various applications, including tasks such as hyper-parameter search [6], neural architecture search [8], continual and incremental learning [22], as well as active learning [27, 5].
+Also, in [41], the authors proposed to model the sample selection procedure as a constrained discrete optimization problem.
+Data pruning proves valuable at reducing memory and computational cost in various applications, including tasks such as hyper-parameter search [6], NAS [8], continual and incremental learning [22], as well as active learning [27, 5].
 Other related fields are dataset distillation and data-free knowledge distillation (DFKD).
@@
 In our paper, we explore the process of distilling knowledge from a model trained on a large dataset to a model trained on a pruned subset of the original data.
-Specifically, we focus on self-distillation and present several striking observations that emerge when integrating SD for data pruning.
+We focus on self-distillation and present several striking observations that emerge when integrating SD for data pruning.
 3 Method
@@
 CIFAR-100 comprises 100 classes, organized into 20 super-classes, each containing 5 sub-classes.
-For example, the super-class "People" contains the classes: "Baby", "Boy", "Girl", "Man", and "Woman".
+For example, the super-class ”People” contains the classes: ”Baby”, ”Boy”, ”Girl”, ”Man”, and ”Woman”.
 As shown in Fig.
-2 (top), the teacher accurately predicts the ground-truth class "Girl" (class index 35) with high confidence while also assigning high confidence values to the classes "Woman" (98), "Man" (46), and "Boy" (11).
+2 (top), the teacher accurately predicts the ground-truth class ”Girl” (class index 35) with high confidence while also assigning high confidence values to the classes ”Woman” (98), ”Man” (46), and ”Boy” (11).
 This ‘dark knowledge’ is valuable for training as it offers a broader view of class hierarchies and data distribution.
@@
 Figure 2: Learning from the teacher predictions.
-An example of soft predictions computed by a teacher model trained on the entire data (top), a model trained on $25\%$ of the data (middle), and a student model trained on $25\%$ of the data with KD (bottom), for an evaluation sample of class "Girl" from CIFAR-100.
+An example of soft predictions computed by a teacher model trained on the entire data (top), a model trained on $25\%$ of the data (middle), and a student model trained on $25\%$ of the data with KD (bottom), for an evaluation sample of class ”Girl” from CIFAR-100.
 Using KD, the student can better learn close or ambiguous categories by leveraging knowledge captured by the teacher from the full dataset.
@@
 4.2.
-Refer to caption (a) CIFAR-100 highest score pruning samples Refer to caption (b) SVHN highest score pruning samples
+(a) CIFAR-100 highest score pruning samples
+(b) SVHN highest score pruning samples
 Figure 3: Highest scoring samples.
-Top 10 highest scoring samples selected by the ‘forgetting’ method for CIFAR-100 and SVHN datasets.
-The labels of the majority of the images are ambiguous or hard to resolve due to class complexity or low image quality.
+Top 10 highest scoring samples selected by the ‘forgetting’ pruning method for CIFAR-100 and SVHN datasets.
+The labels of the majority of the images are ambiguous due to class complexity or low image quality.
+(a) CIFAR-100
+(b) SVHN
+(c) CIFAR-10
+Figure 4: Data pruning results with knowledge distillation.
+Accuracy results across different pruning factors $f$, and various pruning approaches (’forgetting’, EL2N, GraNd and random pruning) on the CIFAR-100, SVHN, and CIFAR-10 datasets.
+We use an equalized weight in the loss (i.e., $\alpha=0.5$).
+Using KD, significant improvement is achieved across all pruning regimes and all pruning methods.
+Random pruning outperforms other pruning methods for low pruning factors.
+For sufficiently high $f$, the accuracy is robust to the choice of the pruning approach in the presence of KD.
 3.3 Theoretical motivation
@@
 Then, we also investigate how the KD weight, the teacher size and the KD method affect student performance under different pruning regimes.
-(a) CIFAR-100
-(b) SVHN
-(c) CIFAR-10
-Figure 4: Data pruning results with knowledge distillation.
-Accuracy results across different pruning factors $f$, and various pruning approaches (’forgetting’, EL2N, GraNd and random pruning) on the CIFAR-100, SVHN, and CIFAR-10 datasets.
-We use an equalized weight in the loss (i.e., $\alpha=0.5$).
-Using KD, significant improvement is achieved across all pruning regimes and all pruning methods.
-Random pruning outperforms other pruning methods for low pruning factors.
-For sufficiently high $f$, the accuracy is robust to the choice of the pruning approach in the presence of KD.
 (a) ImageNet, Top-1 accuracy
@@
 We also utilize a class-balanced random pruning scheme, which, given a pruning budget, randomly and equally draws samples from each class.
-Implementation Details.
-For computational efficiency we conduct our self-distillation experiments on all datasets using the ResNet-32 [13] architecture, except for ImageNet for which we utilize the larger ResNet-50.
-Our training and distillation recipes are simple.
-We utilize SGD with Momentum to optimize the models and incorporate basic data-augmentations during training.
-Additional implementation details can be found in the supplementary.
 4.1 Training on pruned data with KD
@@
 In fact, even on a large scale dataset as ImageNet, comparable accuracy can be achieved by randomly retaining just 30% of the data, while training on larger subsets remarkably results in superior accuracy to the baseline (e.g., +1.6% using a random subset of 70%).
-In contrast, we observe that in the absence of KD, preserving accuracy is only possible in the presence of more data (i.e., more relaxed pruning levels).
-In particular, on CIFAR-100 we found accuracy preservation to be unattainable without KD even in the least aggressive pruning scenario we have tested ($f=0.9$), and regardless of the pruning method.
 Figure 6: Optimal KD weight versus pruning factor.
@@
 Similar results on SVHN can be found in the supplementary.
+Method 5% 10% 30% 50%
+w/o KD 14.46 22.21 49.41 67.47
+KD [15] 28.62 46.27 66.82 70.95
+FitNets [33] 25.66 44.84 65.7 70.77
+AB [14] 30.5 47.68 66.15 71.22
+AT [47] 28.26 42.59 65.75 70.45
+FT [19] 28.34 44.01 64.95 70.75
+FSP [42] 27.62 37.16 62.79 69.72
+NST [17] 26.2 44.5 64.93 70.97
+PKT [31] 27.3 44.09 65.22 70.7
+RKD [30] 21.69 43.03 65.43 70.36
+SP [39] 29.09 42.53 65.62 70.72
+VID [1] 32.5 49.46 67.38 71.16
+Table 1: Comparison of different KD approaches on several pruning levels of CIFAR-100.
+We add various KD loss terms to Eq.
+2, in addition to the vanilla KD term.
+‘Forgetting’ is utilized as the pruning method.
+As observed, integrating VID [1] further improves training on the pruned dataset.
 (a) CIFAR-100. Student architecture: ResNet-32 (RN32).
@@
 We use random pruning with different fractions.
-Surprisingly, under low pruning factors, increasing the teacher’s capacity results in lower student accuracy.
+Interestingly, under low pruning factors, increasing the teacher’s capacity results in lower student accuracy.
 4.3 Using teachers of different capacities
@@
 In Fig.
-6(a), we present accuracy results across various pruning factors for the case of randomly pruning CIFAR-100 and training with a ResNet-32 student.
+7(a), we present accuracy results across various pruning factors for the case of randomly pruning CIFAR-100 and training with a ResNet-32 student.
 We employ 6 teacher architectures of increasing capacities: (1) ResNet-14 with $69.9\%$ accuracy, (2) ResNet-20 with $70.23\%$ accuracy, (3) ResNet-32 with $71.6\%$ accuracy, (4) ResNet-56 with $72.7\%$ accuracy, (5) ResNet-110 with $74.4\%$ accuracy, and (6) WRN-40-2 with $75.9\%$ accuracy.
@@
 Similarly, in Fig.
-6(b) we present results for the same experiment using a ResNet-20 student, while Fig.
-6(c) depicts results of a similar experiment on CIFAR-10 for the ResNet-32 student.
+7(b) we present results for the same experiment using a ResNet-20 student, while Fig.
+7(c) depicts results of a similar experiment on CIFAR-10 for the ResNet-32 student.
 As observed, at low pruning factors, increasing the teacher’s capacity harms the accuracy of the student.
-Conversely, at high pruning factors (e.g.
-$0.9$ and $1$) larger teachers generally lead to improved results, as expected in common KD for model compression.
 This trend is consistently observed across various student architectures and datasets, and is robust to the selection of the KD temperature.
@@
 This observation highlights a striking phenomenon: the capacity gap problem, which denotes the disparity in architecture size between the teacher and student, becomes more pronounced when applying knowledge distillation during training on pruned data.
-Surprisingly, for small pruning fractions, teachers with a smaller capacity than the student can lead to higher student accuracy levels compared to larger teachers.
-The results also suggest that the effectiveness of smaller teachers in improving the student is bounded to an extent by their own accuracy levels.
 4.4 Comparing different KD approaches
-Method 5% 10% 30% 50%
-w/o KD
-14.46
-22.21
-49.41
-67.47
-KD [15]
-28.62
-46.27
-66.82
-70.95
-FitNets [33]
-25.66
-44.84
-65.7
-70.77
-AB [14]
-30.5
-47.68
-66.15
-71.22
-AT [46]
-28.26
-42.59
-65.75
-70.45
-FT [19]
-28.34
-44.01
-64.95
-70.75
-FSP [41]
-27.62
-37.16
-62.79
-69.72
-NST [17]
-26.2
-44.5
-64.93
-70.97
-PKT [31]
-27.3
-44.09
-65.22
-70.7
-RKD [30]
-21.69
-43.03
-65.43
-70.36
-SP [39]
-29.09
-42.53
-65.62
-70.72
-VID [1] 32.5 49.46 67.38 71.16
-Table 1: Comparison of different KD approaches on several pruning levels of CIFAR-100.
-We add various KD loss terms to Eq.
-2, in addition to the vanilla KD term.
-‘Forgetting’ is utilized as the pruning method.
-As observed, integrating VID [1] further improves training on the pruned dataset.
 So far, we have utilized solely vanilla KD during training.
@@
 Finally, for small pruning fractions, we made the surprising observation that the student benefits more from teachers with equal or even smaller capacities than that of its own, over teachers with larger capacities.
-We believe that these empirical observations can pave the way for in-depth research, both theoretical and experimental, to enhance our understanding of the optimization process when training models on pruned data using KD.